OpenStudy (anonymous):

integrate sqrt(1 + sqrt(x)) dx

hartnn (hartnn):

did u try substituting 1+sqrt x = t^2 ??

OpenStudy (anonymous):

y t^2 ?

hartnn (hartnn):

because i wanted to eliminate the sqrt sign....

hartnn (hartnn):

outer sqrt.

hartnn (hartnn):

try that substitution, if u don't get it, i'll help you....

OpenStudy (anonymous):

just a substituion of 1+sqrt x to t would work/

hartnn (hartnn):

yes, that too will work....it will just involve fractional powers.......

OpenStudy (anonymous):

yo yo.............

OpenStudy (anonymous):

I mean you're saying the same thing: Either you do: \[\sqrt{1+\sqrt x}=\psi \implies \psi^2 = 1+\sqrt x\] They are literally the same thing.